OpenStudy (anonymous):

Is the cost of equity calculated from the CAPM model, pre -tax or post-tax? I think it is post-tax but cannot come up with a proper rationale. Can anyone help?

OpenStudy (anonymous):

The existence of pre-tax cost of debt and post-tax cost of debt is due to the acknoledgement of the tax benefit from issuing debt.There is no tax benefit from paying divdends,so it makes no sense talking about pre-tax,post-tax cost of equity for a firm.When you think about cash flow to equity you can only assume that the taxes owed by the company have already been paid.Now, the taxation over the income of the shareholder is a whole different issue that does not take place in this discussion,since it is not taken in consideration either in cost of equity or cost of debt.

OpenStudy (anonymous):

A small enhancement: The cost of equity is usually stated in the form of after-corporate-tax. As stated by Fsbtista1, there is no tax savings derived from the payment of a dividend. However, it is possible to state the cost of equity on a pre-tax basis as well. \[r _{equity, pre-tax}=r _{equity, after-tax}\div \left( 1-t _{corporate} \right)\] This pre-tax cost of equity would be comparable to the pre-tax cost of debt (approximately the interest rate on debt). Either way, what is important to remember is to look at both costs (debt and equity) in a way that is consistent. Both pretax, or both post-tax. The convention is both post-tax.

OpenStudy (anonymous):

A final thought - Beta in the CAPM is pretty much a unitless number. However, when you use the actual SML equation to estimate a cost of capital (i.e. \[E \left[r_{i} \right]=r_{f}+\beta_{i}\left( r _{m}-r_{f} \right)\] you are using the market risk premium and the risk free returns to come up with a required return on equity. The market risk premium is based on stock market returns, including dividend returns. The dividend return is clearly aftertax, and the capital gains are also aftertax. Therefore, the estimated cost of equity using the SML (CAPM) should be after-tax too. I don't know if the risk-free rate is after-tax or not. I suspect that it is a moot point since the government cannot deduct it's interest expense for tax purposes (the government does not apply a tax to itself based on its own P&L statement: it does not really tax its own profits), but by the same reasoning, it is probably closer in spirit to an after tax rate.

OpenStudy (anonymous):

since CAPM is used to estimate equity discounting post-corporate tax, and Beta is calculated using stock returns which are also obviously post-corporate tax, and RFR is really only a PERSONAL opportunity cost, so it makes sense that CAPM has to be post tax.

OpenStudy (anonymous):

It is post corporate tax but pre-personal tax.